### v1

# Measuring Hawking Radiation from Black Hole Morsels in Astrophysical Black Hole Mergers

Abstract
We show that it is possible to observe the Hawking radiation emitted by small black holes assumed to form in catastrophic astrophysical events such as black hole mergers. Gamma ray bursts in the TeV range are unique footprints of these asteroid-mass black hole morsels ejected during the merger. The time delay of the gamma ray bursts from the gravitational wave event is correlated to the mass distribution of the morsels. The integrated mass of the morsels allowed by the unaccounted merger mass leads to a Hawking induced radiation in photons that is above the sensitivity of atmospheric Cherenkov telescopes such as HESS, LHAASO and HAWC.
Black Holes (BHs) are extreme objects predicted by Einstein’s theory of general relativity. Their presence in the Universe has been detected indirectly via their gravitational interactions. We now know that super-massive Black Holes (SMBHs) are present in the centre of most galaxies [1, 2], with masses between $10^{5}$ and several billion solar masses. They reveal their presence by affecting the rotation speed of nearby stars [3, 4] or producing active galactic nuclei. Recently, the first direct images of the region near a SMBH have been taken [5]. Stellar mass BHs are produced at the end of heavy star lifetimes, and they are at the origin of gravitational wave signals observed at LIGO/VIRGO interferometers [6]. A distinctive signal from BHs may be due, theoretically, to the emission of particles via the evaporation mechanism introduced by S.Hawking [7]. However, since their effective temperature is inversely proportional to their mass, astrophysical BHs heavier than a few solar masses are stable because they are colder than the cosmic microwave background. Smaller BHs, however, emit radiation that can be potentially observed. These asteroid-mass BHs could have been produced in the early Universe [8], and constitute a potential candidate for dark matter if sufficiently long-lived [9, 10, 11, 12]. The Hawking radiation stemming from these primordial BHs is constrained by measuring the diffuse gamma ray background [13].
In this letter, we explore the observational consequences of the production of a large number of small BH morsels during a catastrophic event such as the merger of two astrophysical BHs. As we shall show, the Hawking radiation stemming from these BH morsels gives rise to gamma ray bursts (GRBs) possessing a distinctive fingerprint.
Although BHs lighter than a few solar masses are not expected to be produced via gravitational collapse of over-dense stellar matter, one can still envision that, under certain strong non-linearity in the gravitational field within general relativity or beyond, small BHs may form. For an example, primordial BHs are expected to form via quantum space-time fluctuations enhanced by an inflationary period [14, 15, 16]. Gravitational non-linearity does occur in astrophysical events, as shown in [17, 18, 19, 20] for highly rotating BHs, and in [21, 22, 23] for BH merger dynamics. Specifically, non-linearity appears both in merger simulations [22] and during the successive ringdown phase [21, 23]. BH mergers are especially interesting because they can be detected via gravitational waves. Additionally, the events are expected to be followed by GRBs only in the presence of pre-existing material from their progenitor star [24, 25], from supernova remnants [26, 13, 27] or from active galactic nuclei [28]. The emission of small BH morsels provides a novel mechanism for the emission of high-energy particles, within GRBs characterised by the TeV energy scale.
As we remain agnostic about the details of the BH morsel production, we assume the existence of a distribution of masses that escapes the merger. In other words, when the two BHs merge into a single one, regions of space-time characterised by a strong non-linear gravitational field lead to the formation of BH morsels outside the final horizon. The physics of the Hawking radiation of the BH morsels is encoded in their mass distribution function
$\rho(M_{\rm Bm})\equiv\frac{dn_{\rm Bm}}{dM_{\rm Bm}}\,.$ (1)
Here, $\rho_{\rm Bm}\,d{M_{\rm Bm}}$ gives the number of BH morsels in the infinitesimal mass range between $M_{\rm Bm}$ and $M_{\rm Bm}+dM_{\rm Bm}$. For now, we neglect the effect of the gravitational field of the merged BH, and consider the morsels to be in a spherically-symmetric shell configuration. Each morsel emits particles with rates that depend on its mass. The emission rates per each particle species $p$ follow the master equation [29]
$\frac{d^{2}\,N_{p}}{dt\;dE_{p}}=\frac{1}{h}\frac{\Gamma_{p}(E_{p},M_{\rm Bm})}{\exp\frac{E_{p}}{k_{B}T_{\rm Bm}}\pm 1}\,,$ (2)
where the denominator is the usual Boltzmann statistic factor for BH morsel temperature $T_{\text{Bm}}$ (with ±$\pm 1$ for fermions/bosons, respectively) and the numerator contains the specific grey body factors. Hence, the emission rate for the distribution of BH morsels can be computed as
[…]
where $\mu_{t}$ is the mass of the morsels at time $t$ as a function of the initial mass $M_{\text{Bm}}$. The above formula yields the differential primary flux for each particle species $p$. We computed the BH evolution and $J_{p}$ numerically by use of the open-source public code BlackHawk [30, 31], where radiative emission, decays and hadronisation of the primary particles is taken into account by using Pythia [32]. The hadronisation tables, therefore, allow us to reliably generate secondary radiation for primary particle energies between $5$ GeV up to a few TeV. For the numerical results in this letter, we use as template a population of BH morsels with equal masses and non rotating, as angular momentum dissipates faster than the mass of the evaporating BH [33].
For LIGO/VIRGO/KAGRA mergers [34, 35, 36, 37], the BH masses are of stellar order, ranging from a few to several tens of solar masses with associated distance span between 240 Mpc to 3 Gpc. The masses of the initial and final BHs can be indirectly measured from the gravitational wave spectrum, with errors of a few solar masses, depending on the event. For instance, for GW170814 [38], which is the first BH merger observed by all three detectors of LIGO and VIRGO, the initial BH masses have been determined to be $30.5^{+5.7}_{-3.0}$ and $25.3^{+2.8}_{-4.2}$, in units of the solar mass $M_{\odot}$. The final BH mass is $53.2^{+3.2}_{-2.7}\;M_{\odot}$, while the total energy emitted in gravitational waves amounts to $2.7^{+0.4}_{-0.3}\;M_{\odot}$. Taking into account the errors in the parameters, the energy budget of the merger still allows for a total energy of a few solar masses to be emitted in BH morsels. Conservatively, we will take one solar mass as a reference value for the maximal total mass in morsels.
Refer to captionRefer to caption
Figure 1: Left: Differential flux $E_{\gamma}^{2}dF_{\gamma}/dE_{\gamma}$ emitted by a distribution of BH morsels of same mass at a distance of $300$ Mpc. The solid lines correspond to $M_{\text{Bm}}=2\times 10^{7}$ kg at different times: until 500 sec (blue), at 3000 sec (purple) and at 3400 sec (red). The latter is close to the evaporation time. The dotted lines show the primary spectrum. For comparison, in green we show the spectrum for a mass of $10^{9}$ kg, for which the emission remains constant over a long time period. The fluxes are normalised to a total morsel mass of 1 $M_{\odot}$. Right: Evaporation time as a function of the BH mass, computed via BlackHawk.
Via evaporation, the BH morsels emit all kinds of SM particles, however only the neutral stable ones, photons and neutrinos, reach detectors on Earth without being deflected by galactic magnetic fields. To compute template photon and neutrino fluxes, we consider as benchmark a BH merger at a distance $D=300$ Mpc from Earth, corresponding to the range of the closest detected signals. The photon flux on Earth can be computed as
[…]
where we included a geometric factor of $1/2$. In the left panel of Fig. 1 we show the differential fluxes, $E_{\gamma}^{2}dF_{\gamma}/dE_{\gamma}$, for a distribution of BH morsels of same mass, normalised to a total of one $M_{\odot}$. The solid lines provide templates for morsel masses of $2\times 10^{7}$ kg, where the colours indicate different times from the production. As the evaporation lifetime of such mass is about 1 hour, the emission remains constant for timescales of about $500$ sec, emitting the spectrum illustrated by the blue line. Towards the final evaporation time, both the emissivity and the particle energies increase, approaching the ‘explosion’ at the end of the BH lifetime [7]. For the red curve, close to the evaporation time, the primary energies exceed $100$ TeV, reaching the current limit of the BlackHawk code. In fact, the wiggles at high energies correspond to numerical instabilities in the code. We also included the effect of the optical transparency of the intergalactic medium, which cuts off the propagation of photons with energies exceeding $100$ TeV [39], manifesting as a sudden drop in the red curve. For reference, the dotted lines show the primary emission of photons. For different morsel masses, the evaporation time will change drastically as it is proportional to $M_{\text{Bm}}^{3}$. Lighter BHs will, therefore, generate an energetic GRB in a shorter time. Instead, heavier morsels provide a steady signal at the beginning, characterised by reduced particle energies, proportionally to the Hawking temperature $T_{\text{Bm}}\propto M_{\text{Bm}}^{-1}$, while the overall luminosity is reduced. As a reference, we show in dashed green the early curve for a mass of $10^{9}$ kg, corresponding to an evaporation lifetime of about 16 years. Nevertheless, all initial BH morsel masses will generate an energetic GRB with similar characteristics with a time delay from the merger of the order of their evaporation time. In the right panel of Fig. 1 we show the evaporation lifetime as a function of the BH mass, as computed by BlackHawk. Similar fluxes can be obtained for neutrinos.
Multimessenger astrophysics associated to gravitational wave events is part of the LIGO/VIRGO/KAGRA programme, and electromagnetic counterparts are mainly being monitored by Fermi-GBM and Swift-BAT experiments within 30 seconds from all event alerts [40]. Motivated by neutron star mergers, the coverage is between a few keV and MeV. As shown in Fig. 1, we expect very small signals from the BH morsels within this range, well below the current experimental sensitivity of $10^{-7}$ erg sec-1 cm-2 for the flux on Earth. In 2016, however, a short GRB was associated to a BH merger event [26], prompting interest in exploring electromagnetic counterparts of BH mergers [41]. The expected signals are mainly associated with pre-existing material surrounding the BH binary, triggering searches in the optical [42, 43, 44, 45, 46, 47] or keV/MeV [48, 49, 50] ranges. The HESS telescope has followed four BH mergers from the O2 and O3 runs of LIGO/VIRGO [51], covering photon energies between 1 and 10 TeV. However, the observation could only focus on the region of the sky at a time $10^{4}\div 10^{5}$ sec after the BH merger occurred. As no significant point-like source was found in the merger region of the sky, the energy flux from the BH mergers could be constrained to be below $10^{-12}~{}\text{erg}\ \text{sec}^{-1}\ \text{cm}^{-2}$. For morsel masses corresponding to evaporation times within or after the observation window, the HESS data can be used to extract a bound on the total mass that can be emitted as a function of the morsel masses. For each event, we compare the limit on the luminosity to the maximal luminosity of the morsel Hawking radiation within the respective observation window, computed by integrating the source emission between $1$ and $10$ TeV. The results are shown in Fig. 2, where the strongest bounds come from the closest events, GW170814 [38] and GW190728__\__064510. The BH morsel mass ranges stem from the different observation times, where masses smaller than the start of the curves are unconstrained as total evaporation occurs before the HESS observation window. For morsel masses above $15\times 10^{7}$ kg, the energies of the emitted particles fall below the sensitivity range of HESS, and the constraint quickly worsens. This should be taken as a simple estimate, as the limits in [51] were obtained for a $E^{-2}$ power-law spectrum of the photons, which does not exactly match the spectra in Fig. 1.
Figure 2: Estimated upper limit on the total mass $M_{\text{tot}}$ emitted in BH morsels as a function of the morsel mass from the HESS observation of four LIGO/VIRGO events. The curves are extracted from GW170814 (blue), GW190512__\__180714 (red), GW190728__\__064510 (orange) and GW200224 (purple). The upper labels indicate the evaporation time as a function of the BH morsel masses.
As visible signals from BH evaporation always entail photons above the TeV energy, this signal offers a golden opportunity for high energy atmospheric Cherenkov telescopes like HESS, HAWC [52] and LHAASO [53, 54, 55]. The latter recently published a catalogue of multi-TeV sources, reporting the observation of differential fluxes within the range $10^{-13}\div 10^{-11}~{}\text{erg}\ \text{sec}^{-1}\ \text{cm}^{-2}$ [54]. Note that, if the BH morsels have masses above $10^{9}$ kg and evaporation times above several years, BH mergers predating the first LIGO/VIRGO detection would leave behind a relic BH morsel shell and produce a TeV scale GRB uncorrelated to merger catalogues, similar to the signal studied in [56].
Neutrino observatories like ANTARES and IceCube have also monitored BH mergers in search for neutrino counterparts [57]. Both ANTARES [58, 59] and IceCube [60] search for neutrino events within a window of ±$\pm 500$ sec around the BH merger, finding no excess over the expected backgrounds. However, the limits in terms of fluxes are orders of magnitude above the ones for photons, requiring typical source luminosity of $10^{51}$ erg sec-1. As the BH morsels predict similar fluxes in photons and neutrinos, such bounds are not competitive.
Several SMBH binary systems are known in galaxies, probably following the merger of two progenitor galaxies which hosted a SMBH at their centre. It is believed that the galaxy merger generates a binary system at distances down to 1 pc, while further decay leading to a merger is due to the interactions with the star population [61]. Such merger would produce gravitational waves in the sensitivity ranges of Pulsar timing array telescopes (PTA) [62] or LISA. Intriguingly, it has recently been reported that the activity of the galaxy SDSS J1430+2303, at a distance of $378$ Mpc from Earth hints at a possible SMBH merger within three years [63, 64]. The system is currently being monitored by HESS [65], while a gravitational wave signal will be in the observation window of PTAs.
Because of the much larger masses, a SMBH merger can potentially emit much more mass in BH morsels, hence providing more intense GRBs than a stellar mass merger. Furthermore, the initial photons emitted by heavier morsels could be visible and provide steady signals that would outlive the photons emitted by other astrophysical sources. Another interesting feature is the fact that the BH morsels are produced and evaporate within the strong gravitational field of the merged SMBH, hence their spectrum may record a signature of where the BH morsels have been produced and the mass of the merged SMBH. For illustration purposes, and for simplicity, we computed the gravitational effects considering a Schwarzschild SMBH of mass $10^{8}\;M_{\odot}$ and a population of morsels of mass $10^{9}$ kg produced at a given distance and energy. We include the main effects due to the time dilation and energy redshift induced by the SMBH gravity. In Fig. 3, we show the expected differential flux at different times for BH morsels produced at a distance of $0.5$ Schwarzschild radius from the SMBH horizon. At early times, the flux is reduced and redshifted, while it approaches the asymptotic spectrum (blue curve) after about $50.000$ sec. After that, the emission remains constant until the evaporation time is reached. The timing of the effect depends crucially on the initial energy of the BH morsels at production.
Figure 3: Differential flux on Earth from one solar mass of BH morsels of individual masses $10^{9}$kg at a distance of $300$ Mpc. We assume that the morsels were produced at a distance $0.5$ Schwarzschild radius from the horizon of the merged SMBH, and with energy $\frac{E}{M_{\text{Bm}}c^{2}}=1+10^{-7}$. The spectrum is heavily distorted at initial times by the gravitational field of the merged SMBH of mass $10^{8}\ M_{\odot}$.
So far, we have presented simple templates based on BH morsels of the same mass emitted radially as a spherically symmetric shell. An extended mass distribution can modulate and prolong the duration of the TeV photon emission. The geometry of the emission, instead, has a smaller impact on the signal, as BHs emit particles in a spherically symmetric manner. Hence, as long as the morsels are not screened by the merged BH, their Hawking radiation will reach our detectors. Some distortions of the spectrum can be expected if the BH morsels travel with relativistic velocity, in which case their direction of motion with respect to the line of sight becomes important. When BH morsels are emitted in a strong gravitational field, such as for SMBH mergers, their Hawking emission may be affected and a dedicated study is required to quantify such effects, which we leave for future investigations. As a final remark, we recall that BH morsels will most intensely emit at times close to their evaporation and with primary energies exceeding the multi-TeV scale. Besides a correct characterisation of the emission of secondary spectra [66], the Hawking radiation at such energies can be also impacted by the limited validity of the standard model of particle physics, and some new physics effects may induce measurable distortions. The observation of Hawking radiation from BH morsels, therefore, could enlighten us not only about the production of such morsels, but also about particle physics at energies beyond the reach of current and future collider experiments, carrying imprints from new physics based on supersymmetry, composite dynamics, or extra dimensions, to name a few.
In summary, in this letter we have shown that it is possible to observe the Hawking radiation emitted by BH morsels assumed to form in catastrophic astrophysical events such as black hole mergers. GRBs are unique footprints of these asteroid-mass black hole morsels emitted after the merger, with photon energies exceeding the TeV scale. The time delay between the gravitational wave event and the GRB is correlated to the mass distribution of the morsels. We showed that the total morsel mass allowed by the unaccounted merger mass leads to a Hawking induced radiation in photons that is above the sensitivity of atmospheric Cherenkov telescopes such as HESS, LHAASO and HAWC.
Acknowledgements
We thank Alexandre Arbey for help and assistance with BlackHawk. We also thank Zhi-Wei Wang and Juri Smirnov for inspiring discussion at the beginning of this project. The work of F.S. is partially supported by the Carlsberg Foundation, semper ardens grant CF22-0922.